# Universal Scaling Solution for a Rigidity Transition: Renormalization Group Flows Near the Upper Critical Dimension

### v1

@@ Abstract
Rigidity transitions induced by the formation of system-spanning disordered rigid clusters, like the jamming transition, can be well-described in most physically relevant dimensions by mean-field theories. However, we lack a theoretical description of these kinds of rigidity transitions that can detect an upper critical dimension below which mean-field theory fails and predict corrections to the theory in and below this upper critical dimension. Towards that end, we show that the critical exponents for a continuous isotropic rigidity transition predicted by a simple dynamical effective medium theory (the Coherent Potential Approximation, CPA) are not mean-field-like, dimension-independent quantities. Instead, we find that the theory has an upper critical dimension $d_{\mathrm{upper}}=2$, in which there are the usual logarithmic corrections and below which the critical exponents are modified. We recapitulate Ken Wilson’s phenomenology of the $(4-\epsilon)$-dimensional Ising model, but with the upper critical dimension reduced to $2$. We interpret this using normal form theory as a transcritical bifurcation in the RG flows, and posit their explicit forms, incorporating a dangerously irrelevant variable becoming marginal in $d=2$. We also derive universal scaling functions from the CPA sufficient to predict all linear response in randomly diluted isotropic elastic systems. These results indicate the types of corrections one expects to disordered rigidity transitions in $2$ dimensions and provide a pathway to understanding the critical phenomena in such systems.
@@ I Introduction
Our model is a continuum version of the coherent potential approximation (CPA) [15, 12, 16]. It is an outgrowth of what is termed rigidity percolation (RP) [17, 18, 19], where a network of springs connecting nodes with no angular forces is diluted, not until it becomes completely disconnected (percolation) but until its elastic moduli vanish. It is known that rigidity percolation on a two-dimensional triangular lattice has critical exponents that differ from those found for rigidity percolation on graphs generated from jammed packings of spheres, with a modulus that with a higher power with excess coordination number [20, 21]. We conjecture that lattice anisotropy [22], and/or undeformed springs in a line (forming second-order constraints [19, 23]) are relevant perturbations at the rigidity transition for spring networks, and that our theory is applicable to a randomly diluted isotropic network without these buckling transitions (Section V). Our continuum theory, when used to describe jamming [12, 24], matches most of the properties seen numerically, both directly and in spring network models generated from jammed configurations [25]. The calculation presented here focuses on the case where the bulk modulus does not jump, which includes most of the rigidity transitions other than jamming. Nevertheless, this calculation captures the predicted scaling of the shear mode close to the jamming transition [24].
@@
Our work is similar in some aspects to that of a preprint by Vogel et al. [41]. They form a self-consistent theory for the shear response of a nearly unjammed solid with similar structure to our continuum CPA. Their analysis incorporates nontrivial momentum dependence which is important in the disordered glassy phase studied by the authors of the article in previous work [42, 43] and by others [44], which is ignored by the CPA. Our work, on the other hand, is focused on calculating the universal scaling functions for the transition, deducing nontrivial normal forms for renormalization group flows, and understanding specifically the singular behavior of the theory found in two dimensions.
The organizational structure of our paper is as follows: in Section II, we briefly review the CPA as applied to weakened elastic media. We show that, under quite general assumptions, elastic moduli vanish linearly in deviations from the critical dilution fraction $\mu\sim\delta p\sim\delta z$. In Section III, we evaluate a dynamical version of the theory near the critical point directly in $d=3$ dimensions and in $d=2$ dimensions, and show that the appropriate scaling for the dynamical behavior close to the critical point is different in these two cases. In Section IV and Appendix C, we cast solutions clost to the critical point into a scaling theory in general dimensions $d$, and show that $d=2$ demarcates the boundary between two differing sets of critical exponents. We construct renormalization group flow equations that are consistent with the critical exponents predicted by the theory (with more details located in Appendices A and B), with an exchange of stability between two RG fixed points in $d=2$. In this way, we interpret $d=2$ as the upper critical dimension of the transition, and find the appropriate scaling variables in $d=2$ analogously to those in the 4D Ising model. The logarithmic shifts that are one signature of the upper critical dimension are investigated in more detail in Appendix D through a direct comparison to a lattice CPA calculation on the bond-diluted triangular lattice. Finally, in Section V, we discuss the applicability of this model more broadly to a wide variety of different rigidity transitions in disordered elastic systems.
@@ II The CPA and critical exponents for static moduli
We examine a continuum version of the coherent potential approximation (CPA) inspired by the lattice CPA [15]. The lattice CPA does not accurately describe the critical properties of the corresponding lattice RP transitions, but they do share critical exponents with wide classes of other rigidity transitions – including the rigidity percolation of statistically isotropic spring networks generated from jammed packings. The lattice CPA can be used to describe a system comprised of purely harmonic springs of strength $k_{0}$ on some regular lattice that are independently randomly occupied with probability $p$. This is equivalent to placing a probability distribution on the strengths of bonds $k^{\prime}$
@@
where the integral is an average over the first Brillouin zone, $\widetilde{z}$ denotes the average number of constraints per microscopic unit in the undiluted system (in the lattice case, it is the number of bonds per node), $\mu_{\textsc{F}}$ is the shear modulus in the completely filled, non-disordered system, and D and G are the dynamical matrix and Green’s function for the medium, respectively. This expression is self-consistently solved for $\mu\left(p,\omega\right)$, which is proportional to the best effective value of the microscopic stiffness $k$.
@@ III.1 Scaling in 3 dimensions
This expansion is more general and also works for the overdamped case where $w\sim i\gamma\omega$. Now we are prepared to define more invariant scaling variables and our first universal scaling function. We first insert the definition of $M$ in for $\mu$ everywhere. Then, we define $\Omega^{2}$ (named for the undamped case) as the scaling variable for $w$ which makes the dominant term of the scaling for the frequency-dependent part $\left|\delta p\right|\Omega^{2}/M$. The definition of $w$ in terms of $\Omega$ is then inserted in the higher-order frequency piece, and the remaining terms are all absorbed into a scaling combination $G$ for an irrelevant variable. The result, after dividing both sides by $\left|\delta p\right|$ and throwing away the higher-order contribution from the denominator of the left-hand side, is
@@
The sign ±$\pm 1=\delta p/\left|\delta p\right|$ is for the rigid and floppy side of the transition, respectively; formulas for $\omega_{0}$ and $g/g_{0}$ can be found in Appendix C. This is an implicit definition of a universal scaling function for $M=\mathcal{M}\left(\Omega,G\right)$. [^1] In a more complete self-consistent theory beyond the CPA such as [41], the shear modulus will also have a $q$-dependence, leading to a scaling form $\mathcal{M}\left(\Omega,Q,G\right)$. Setting the leading irrelevant piece $G$ to 00 allows us to solve a quadratic equation for $M$ as $2\mathcal{M}\left(\Omega,0\right)=\pm 1-\sqrt{1-4\Omega^{2}}$, as found indirectly in [12] and directly in [16, 38]. However, this form of the universal scaling function (unphysically) has no scattering-induced dissipation on the rigid side of the transition until $\Omega=1/2$, while the full solution retaining $G$ has $\textrm{Im}\left(\mathcal{M}\right)<0$ for all $\Omega>0$. As mentioned in the introduction to this section, this identifies $G$ as an invariant scaling combination associated with a dangerously irrelevant variable for the case of undamped microscopic dynamics, as the function $\mathcal{M}(\Omega,G)$ is not analytic in its second argument at zero. Retaining $G$ is necessary to understand the details of the low-frequency viscosity, susceptibility, and density of states.
For models with many soft modes, we can illustrate the excess number of soft modes by comparing the density of states with that of the Debye model for a crystal, in which $D\left(\omega\right)\sim\omega^{d-1}$. We plot the predicted universal scaling forms for the density of states divided by the Debye form for different values of the dangerously irrelevant variable in Figure 2. The peak in the excess density of states is located near the frequency $\omega^{*}\sim\left|\delta p\right|$ where the density of states becomes nearly flat: a characteristic feature of all rigidity transitions in this family.
Figure 2: Prediction for the shape of the excess density of states near the rigid-floppy transition in $d=3$. We compare our density of states to the Debye result and find an excess of states that contribute to the boson peak that is often seen in disordered rigid systems. The dangerously irrelevant variable $G$ must be retained to capture the Debye phonon contribution to the density of states below $\Omega=1/2$. These long-wavelength phonons are of course important to the physics, but are swamped near the rigidity transition by the flat density of states above $\omega^{*}$. Hence the phonon contribution is irrelevant in the RG sense, even though it is important to the physics.
@@ III.2 Scaling in 2 dimensions
We are interested in the low-frequency behavior. There is now quite clearly a logarithmic singularity at low frequencies. Keeping only the leading-order low frequency terms, we have
@@ IV.3 Scaling variables and scaling functions
Is this interesting, given that we typically don’t do simulations of colloidal gels or other disordered rigid systems below two dimensions? While we do not expect experiments to test our predictions in non-integer dimensions, in Section V we discuss applications to theoretical simulations in $d<2$. In two dimensions, however, these predictions are in principle experimentally observable. This is a case where the upper critical dimension of a theory is physically relevant (at least as determined within the CPA). In the Ising model at its upper critical dimension, there is a logarithmic singularity in the magnetization, which also has important (very slowly vanishing at the critical point) log-log corrections. Here, the situation is identical. Suppose one picks out a physically relevant piece of dynamical information at this rigidity transition, like the frequency $\omega^{*}$ where we cross over into being dominated by dissipation $\textrm{Re}\left(\mu\left(\omega^{*}\right)\right)=-\textrm{Im}\left(\mu\left(\omega^{*}\right)\right)$. For all $d$ away from $2$, we can write how this physical frequency scales with our excess contact number: in the undamped case, $\omega^{*}\sim\delta p^{(4-g_{c})/(4-2g_{c})}$. But in 2 dimensions, there are detectable shifts in this frequency, and it turns out that for the undamped case $\omega^{*}\sim\left|\delta p\right|\left|\log\left|\delta p\right|\right|^{-1/2}$ with additional important log-log corrections.
@@ V Applicability of the transition to physical systems
To what systems do we expect our continuum CPA will give quantitatively accurate predictions? What features of our continuum CPA do we expect to apply more broadly to rigidity transitions? Here we discuss clearly where our theory does not apply, and speculate about where it may provide qualitative or quantitative guidance.
@@
As noted in the introduction and in Appendix D, our continuum CPA does not correctly describe the diluted triangular spring lattice, which exhibits non mean-field critical exponents in its static properties in two dimensions rather than the predicted log-corrections to mean-field theory [20, 21]. Our continuum theory, generalized to describe the abrupt jump in the bulk modulus, applies qualitatively to simulations of spring network geometries generated from jammed packings [49]. We reproduce the numerically observed mean-field exponents in three dimensions, and both this calculation and the simulation find mean-field exponents with log corrections to scaling in two dimensions (however, see below). Indeed, based partly on a discussion years ago with Carl Goodrich, we conjecture that randomly diluting a spring network generated from a jammed packing starting in the rigid phase will undergo a transition where both bulk and shear moduli grow continuously, described quantitatively by the version of our continuum CPA analyzed here.
@@
Why is the triangular lattice (and, by implication, many other spring lattices) not behaving according to our theory? While the triangular lattice is statistically isotropic in its elastic moduli, many critical points are more sensitive than elastic theory to breaking of rotational invariance. The XY model is unstable to breaking of triangular and square symmetries [50, 51], and diffusion-limited aggregation is famous for breaking rotational invariance in a way that revealed itself only in (then) large simulations [22]. Statistically isotropic lattices might show mean-field behavior. Another likely culprit are the straight lines between bonds in the triangular lattice, which are shared with the Mikado networks [52] formed by random long fibers cross-linked at their intersections. A node connecting two parallel bonds cannot move freely under tensile stress, while at any non-zero angle it needs a third constraint to fix it in place: such second-order constraints are expected to change critical properties [19, 23]. A generic lattice with the same connectivity structure as a regular triangular lattice, as suggested by Jacobs [19], could be enough to show mean-field behavior.
@@
As noted above, this continuum analysis has a natural extension to describe a jamming transition, inspired again by [12, 16, 38]. There the upper critical dimension is believed to be two, and our model (and the lattice CPA models) does agree with known critical exponents. However, our calculation for jamming shows logarithmic corrections to scaling that only arise in frequency-dependent quantities, as we have seen in this manuscript, while convincing numerical work shows logarithmic corrections to the finite-size scaling of zero-frequency elastic moduli in jamming simulations [49]. We suspect that the lack of logarithmic corrections to the finite-size effects in our 2D jamming variant is related to the CPA assumption that the moduli are independent of wavevector. One notes that the Green’s function and other wavevector-dependent properties have a wavevector scale in 2D that does have logarithmic corrections to scaling $q^{*}\sim\left|\delta p\right|^{1/2}\left|\log\left|\delta p\right|\right|^{-1/2}$ (similar to the frequency scaling). Our current theory replaces the system near its rigidity transition with a uniform but frequency dependent modulus, ignoring the important effects of spatial fluctuations. A more sophisticated self-consistent approach allowing for more general forms of response functions, such as the preprint of Vogel et al. [41] mentioned in the Introduction, could be analyzed to extract a universal scaling theory that not only correctly describes the finite-size effects in two-dimensional jamming, but also describes the 2D and 3D crossovers to the elastic-dipole induced correlations of glasses [42, 43, 53] and Rayleigh scattering.

### v2

@@ Abstract
Rigidity transitions induced by the formation of system-spanning disordered rigid clusters, like the jamming transition, can be well-described in most physically relevant dimensions by mean-field theories. A dynamical mean-field theory commonly used to study these transitions, the coherent potential approximation (CPA), shows logarithmic corrections in $2$ dimensions. By solving the theory in arbitrary dimensions and extracting the universal scaling predictions, we show that these logarithmic corrections are a symptom of an upper critical dimension $d_{\mathrm{upper}}=2$, below which the critical exponents are modified. We recapitulate Ken Wilson’s phenomenology of the $(4-\epsilon)$-dimensional Ising model, but with the upper critical dimension reduced to $2$. We interpret this using normal form theory as a transcritical bifurcation in the RG flows and extract the universal nonlinear coefficients to make explicit predictions for the behavior near $2$ dimensions. This bifurcation is driven by a variable that is dangerously irrelevant in all dimensions $d>2$ which incorporates the physics of long-wavelength phonons and low-frequency elastic dissipation. We derive universal scaling functions from the CPA sufficient to predict all linear response in randomly diluted isotropic elastic systems in all dimensions.
@@ I Introduction
Our model is an isotropic, continuum version of the coherent potential approximation (CPA) [15, 12, 16]. It replicates the CPA predictions for the dilution of a random, amorphous spring network published by Düring et al. [17]. Both are outgrowths of what is termed rigidity percolation (RP) [18, 19, 20], where a network of springs connecting nodes with no angular forces is diluted, not until it becomes completely disconnected (percolation) but until its elastic moduli vanish. It is known that rigidity percolation on a two-dimensional triangular lattice has critical exponents that differ from those found for rigidity percolation on graphs generated from jammed packings of spheres, with a modulus that with a higher power with excess coordination number [21, 22]. We conjecture that lattice anisotropy [23], and/or undeformed springs in a line (forming second-order constraints [20, 24]) are relevant perturbations at the rigidity transition for spring networks, and that our theory is applicable to a randomly diluted isotropic network without these buckling transitions (Section V). This continuum theory, when used to describe jamming [12, 25], matches most of the properties seen numerically, both directly and in spring network models generated from jammed configurations [26] and simulations of diluted amorphous spring networks [17]. The calculation presented here focuses on the case where the bulk modulus does not jump, which includes most of the rigidity transitions other than jamming. Nevertheless, this calculation captures the predicted scaling of the shear mode close to the jamming transition [25]. We leverage our exact solutions of the continuum theory to generate universal scaling forms for all linear response properties, and we use them to calculate new critical exponents below $2$ dimensions described by nonlinear renormalization group flows. We identify a scaling variable which is dangerously irrelevant for $d>2$ which is responsible for low-frequency dissipation, the phonon density of states, and the logarithmic corrections in $d=2$.
@@
Our work is similar in some aspects to that of a preprint by Vogel et al. [42]. They form a self-consistent theory for the shear response of a nearly unjammed solid with similar structure to our continuum CPA. Their analysis incorporates nontrivial momentum dependence which is important in the disordered glassy phase studied by the authors of the article in previous work [43, 44] and by others [45]; this type of momentum-dependent modulus is ignored by the CPA. Our work, on the other hand, is focused on calculating the universal scaling functions for the transition, deducing nontrivial normal forms for renormalization group flows, and understanding specifically the singular behavior of the theory found in two dimensions.
The organizational structure of our paper is as follows: in Section II, we briefly review the CPA as applied to weakened elastic media. We show that, under quite general assumptions, elastic moduli vanish linearly in deviations from the critical dilution fraction $\mu\sim\delta p\sim\delta z$. In Section III, we evaluate the universal scaling functions for the space-time linear response of the theory near the critical point directly in $d=3$ dimensions and in $d=2$ dimensions, and show that the appropriate scaling for the dynamical behavior close to the critical point has log corrections in two dimensions, as seen earlier [17]. In Section IV and Appendix C, we cast solutions close to the critical point into a scaling theory in general dimensions $d$, and show that $d=2$ demarcates the boundary between two differing sets of critical exponents. We construct renormalization group flow equations that are consistent with the critical exponents predicted by the theory (with more details located in Appendices A and B), with an exchange of stability between two RG fixed points in $d=2$. In this way, we interpret $d=2$ as the upper critical dimension of the transition, and find the appropriate scaling variables in $d=2$ analogously to those in the 4D Ising model. The logarithmic shifts that are one signature of the upper critical dimension are investigated in more detail in Appendix D through a direct comparison to a lattice CPA calculation on the bond-diluted triangular lattice. Finally, in Section V, we discuss the applicability of this model more broadly to a wide variety of different rigidity transitions in disordered elastic systems.
@@ II The CPA and critical exponents for static moduli
We examine a continuum version of the coherent potential approximation (CPA) inspired by the lattice CPA [15]. The lattice CPA can be used to describe a system comprised of purely harmonic springs of strength $k_{0}$ on some regular lattice that are independently randomly occupied with probability $p$. This is equivalent to placing a probability distribution on the strengths of bonds $k^{\prime}$
@@
where the integral is an average over the first Brillouin zone, $\widetilde{z}$ denotes the average number of constraints per microscopic unit in the undiluted system (in the lattice case, it is the number of bonds per node), $\mu_{\textsc{F}}$ is the shear modulus in the completely filled, non-disordered system, and D and G are the dynamical matrix and Green’s function for the medium, respectively. This long-wavelength limit of the lattice problem reproduces the isotropic CPA for amorphous spring networks investigated by Düring et al. in [17]. This expression is self-consistently solved for $\mu\left(p,\omega\right)$, which is proportional to the best effective value of the microscopic stiffness $k$.
@@ III.1 Scaling in 3 dimensions
This expansion is more general and also works for the overdamped case where $w\sim i\gamma\omega$. This asymptotic analysis of the leading correction to the CPA result reproduces the calculation done in Appendix B of [48]. Now we are prepared to define more invariant scaling variables and our first universal scaling function. We first insert the definition of $M$ in for $\mu$ everywhere. Then, we define $\Omega^{2}$ (named for the undamped case) as the scaling variable for $w$ which makes the dominant term of the scaling for the frequency-dependent part $\left|\delta p\right|\Omega^{2}/M$. The definition of $w$ in terms of $\Omega$ is then inserted in the higher-order frequency piece, and the remaining terms are all absorbed into a scaling combination $G$ for an irrelevant variable. The result, after dividing both sides by $\left|\delta p\right|$ and throwing away the higher-order contribution from the denominator of the left-hand side, is
@@
[…]
For models with many soft modes, we can illustrate the excess number of soft modes by comparing the density of states with that of the Debye model for a crystal, in which $D\left(\omega\right)\sim\omega^{d-1}$. We plot the predicted universal scaling forms for the density of states divided by the Debye form for different values of the dangerously irrelevant variable in Figure 2. The peak in the excess density of states is located near the frequency $\omega^{*}\sim\left|\delta p\right|$ where the density of states becomes nearly flat: a characteristic feature of all rigidity transitions in this family. The dangerous irrelevant variable $G$ controls the contiuum phonon density of states, which vanishes in the appropriate scaling limit.
Figure 2: Prediction for the shape of the excess density of states near the rigid-floppy transition in $d=3$ (see also Fig. 5). We compare our density of states to the Debye result and find an excess of states that contribute to the boson peak that is often seen in disordered rigid systems. The dangerously irrelevant variable $G$ must be retained to capture the Debye phonon contribution to the density of states below $\Omega=1/2$. These long-wavelength phonons are of course important to the physics, but are swamped near the rigidity transition by the flat density of states above $\omega^{*}$. Hence the phonon contribution is irrelevant in the RG sense, even though it is important to the physics.
@@ III.2 Scaling in 2 dimensions
We are interested in the low-frequency behavior. There is now quite clearly a logarithmic singularity at low frequencies, as discussed in [17] [^3] Their parameter $\varepsilon$ is our $-\lim_{\omega\rightarrow 0}f$ on the floppy side of the transition with undamped dynamics.. Keeping only the leading-order low frequency terms, we have
@@ IV.3 Scaling variables and scaling functions
Is this interesting, given that we typically don’t do simulations of colloidal gels or other disordered rigid systems below two dimensions? While we do not expect experiments to test our predictions in non-integer dimensions, in Section V we discuss applications to theoretical simulations in $d<2$. In two dimensions, however, these predictions are in principle experimentally observable. This is a case where the upper critical dimension of a theory is physically relevant (at least as determined within the CPA). In the Ising model at its upper critical $d=4$, there is a logarithmic singularity in the magnetization, which also has important (very slowly vanishing at the critical point) log-log corrections. Here, the situation is identical. Suppose one picks out a physically relevant piece of dynamical information at this rigidity transition, like the frequency $\omega^{*}$ where we cross over into being dominated by dissipation $\textrm{Re}\left(\mu\left(\omega^{*}\right)\right)=-\textrm{Im}\left(\mu\left(\omega^{*}\right)\right)$. For all $d$ away from $2$, we can write how this physical frequency scales with our excess contact number: in the undamped case, $\omega^{*}\sim\delta p^{(4-g_{c})/(4-2g_{c})}$. But in 2 dimensions, there are detectable shifts in this frequency, and it turns out that for the undamped case $\omega^{*}\sim\left|\delta p\right|\left|\log\left|\delta p\right|\right|^{-1/2}$ with additional important log-log corrections.
@@ V Applicability of the transition to physical systems
To what systems do we expect the behavior to be quantitatively described by our universal scaling predictions? What features of our predictions do we expect to apply more broadly to rigidity transitions? Here we discuss clearly where our universal predictions do not apply, and speculate about where they may provide qualitative or quantitative guidance.
@@
As noted in the introduction and in Appendix D, our continuum CPA does not correctly describe the diluted triangular spring lattice, which exhibits non mean-field critical exponents in its static properties in two dimensions rather than the predicted log-corrections to mean-field theory [21, 22]. Generalized to describe the abrupt jump in the bulk modulus, the continuum CPA applies qualitatively to simulations of spring network geometries generated from jammed packings [52]. The generalized continuum CPA reproduces the numerically observed jamming mean-field exponents in three dimensions, and both this calculation and the simulation find mean-field exponents with log corrections to scaling in two dimensions (however, see below). Indeed, based partly on a discussion years ago with Carl Goodrich, we conjecture that randomly diluting a spring network generated from a jammed packing starting in the rigid phase will undergo a transition where both bulk and shear moduli grow continuously, described quantitatively by the version of our continuum CPA analyzed here [53].
@@
Why is the triangular lattice (and, by implication, many other spring lattices) not behaving according to our theory? While the triangular lattice is statistically isotropic in its elastic moduli, many critical points are more sensitive than elastic theory to breaking of rotational invariance. The XY model is unstable to breaking of triangular and square symmetries [54, 55], and diffusion-limited aggregation is famous for breaking rotational invariance in a way that revealed itself only in (then) large simulations [23]. Statistically isotropic lattices might show mean-field behavior. Another likely culprit are the straight lines between bonds in the triangular lattice, which are shared with the Mikado networks [56] formed by random long fibers cross-linked at their intersections. A node connecting two parallel bonds cannot move freely under tensile stress, while at any non-zero angle it needs a third constraint to fix it in place: such second-order constraints are expected to change critical properties [20, 24]. A generic lattice with the same connectivity structure as a regular triangular lattice, as suggested by Jacobs [20], could be enough to show mean-field behavior. And diluted, statistically isotropic spring networks with random node positions show mean-field exponents [48].
@@
As noted above, this continuum analysis has a natural extension to describe a jamming transition, inspired again by [12, 16, 39]. There the upper critical dimension is believed to be two, and our model (and the lattice CPA models) does agree with known critical exponents. However, our calculation for jamming shows logarithmic corrections to scaling that only arise in frequency-dependent quantities, as we have seen in this manuscript, while convincing numerical work shows logarithmic corrections to the finite-size scaling of zero-frequency elastic moduli in jamming simulations [52]. We suspect that the lack of logarithmic corrections to the finite-size effects in our 2D jamming variant is related to the CPA assumption that the moduli are independent of wavevector. One notes that the Green’s function and other wavevector-dependent properties have a wavevector scale in 2D that does have logarithmic corrections to scaling $q^{*}\sim\left|\delta p\right|^{1/2}\left|\log\left|\delta p\right|\right|^{-1/2}$ (similar to the frequency scaling). Our current theory replaces the system near its rigidity transition with a uniform but frequency dependent modulus, ignoring the important effects of spatial fluctuations. A more sophisticated self-consistent approach allowing for more general forms of response functions, such as the preprint of Vogel et al. [42] mentioned in the Introduction, could be analyzed to extract a universal scaling theory that not only correctly describes the finite-size effects in two-dimensional jamming, but also describes the 2D and 3D crossovers to the elastic-dipole induced correlations of glasses [43, 44, 57] and Rayleigh scattering. However, the detailed simulations of [17] do appear to reproduce the logarithmic shifts in the density of states and $\ell_{c}^{-1}\sim q^{*}$ explained in this work.